As an Investment Banker your client, Alpha Limited is looking at a restructuring its business. You extract the following data from the financials of the company.



Particulars Rs. Crs.

Equity capital (Face value Rs. 10) 1,000.00

Debentures (@ 12%) 400.00

Long Term unsecured loan (@15%) 200.00

Total 1,600.00


The company has been paying a dividend of 20% per annum (historically). The stock of LMN Limited is listed at Rs. 20 on the NSE.


You are required to


a. Compute the cost of Equity capital

b. Weighted Average cost of capital of Alpha Limited.


Detailed calculation is required



1
Expert's answer
2021-09-29T11:34:57-0400

a. Equity capital:


The cost of equity can be found as:

"Equity\space capital=\frac{interest\space rate\space on\space dividends\times100}{Equity\space capital}=\frac{20\times100}{10 000}=0.2"

Equity capital=0.2


b.

"WACC=keWe+kdWd"


"WACC=keWe+kdWd=0.2\times\frac{10000}{160000}+0.12\times\frac{4000}{160000}+0.15\times\frac{2000}{160000}=0.017" or 1.7%